OpenStudy (anonymous):

If f(x)=3x+2 and g^-1(f^-1(x))=5x-4 find g(x)

OpenStudy (anonymous):

\(\Large f(x)=3x+2 \rightarrow f^{-1}(x)=\frac{x-2}{3} \) now let \(\large y=\frac{x-2}{3} \) so that: \(\Large g^{-1}(f^{-1}(x))=5x-4 \) or \(\Large g^{-1}(y)=5x-4 \) since \(\large x=3y+2 \), \(\Large g^{-1}(y)=5x-4 =5(3y+2)-4=15y+10-4=15y+6\) can you take it from here to find g(x) ??

OpenStudy (anonymous):

g(x)=(x-6)/15??????

OpenStudy (anonymous):

that should be right... yes...

OpenStudy (anonymous):

HECK FREAKING YES!!!!!! THANK YOU SO MUCH.

OpenStudy (anonymous):

yw.... :)